# Parameter estimation of range-migrating targets using OTFS signals from LEO satellites

### v2

@@ Abstract
This study investigates a communication-centric integrated sensing and communication (ISAC) system that utilizes orthogonal time frequency space (OTFS) modulated signals emitted by low Earth orbit (LEO) satellites to estimate the parameters of space targets experiencing range migration, henceforth referred to as high-speed targets. Leveraging the specific signal processing performed by OTFS transceivers, we derive a novel input-output model for the echo generated by a high-speed target in scenarios where ideal and rectangular shaping filters are employed. Our findings reveal that the target response exhibits a sparse structure in the delay-Doppler domain, dependent solely upon the initial range and range-rate; notably, range migration causes a spread in the target response, marking a significant departure from previous studies. Utilizing this signal structure, we propose an approximate implementation of the maximum likelihood estimator for the target’s initial range, range-rate, and amplitude. The estimation process involves obtaining coarse information on the target response using a block orthogonal matching pursuit algorithm, followed by a refinement step using a bank of matched filters focused on a smaller range and range-rate region. Finally, numerical examples are provided to evaluate the estimation performance.
@@ I Introduction
Future wireless networks and Internet-of-Things (IoT) will utilize terrestrial and non-terrestrial stations to ensure seamless and ubiquitous connectivity [1]. Among spaceborne transceivers, low Earth orbit (LEO) satellites are attracting much attention for their lower latency and higher data rates [2]; also, they may efficiently provide integrated sensing and communications (ISAC) functionalities over large regions with the same signal waveform and hardware architecture [3, 4, 5].
@@
Orthogonal time frequency space (OTFS) modulation is a promising transmission scheme for future LEO constellations [6, 7]. OTFS-based transceivers exploit channel diversity across both time and frequency, thus providing better communication performance in doubly-dispersive fading channels as compared to schemes using orthogonal frequency division multiplexing [8, 9, 10]. They are also suitable for radar sensing, as the underlying signal processing resembles that of pulsed-Doppler radars [11]; in particular, the range and range-rate of prospective scattering objects can be inferred from the estimated delay-Doppler channel response [12]. These unique features enable OTFS-based LEO satellites to inherently support ISAC in various forms [13, 14]. Unlike terrestrial stations, LEO-based stations have the capability to also cover aerial and space regions, and this expanded coverage opens up new application scenarios, such as real-time environmental monitoring, enhanced global navigation systems, and space exploration. Communication-centric ISAC solutions appear particularly promising in this context, as existing data signals can be opportunistically reused to add the sensing functions at a low cost, without altering the underlying communication protocol or affecting its performance [12, 15, 16, 17, 18, 19]; this approach is indeed considered in this study.
@@
Several solutions for signal estimation and detection have been deployed for OTFS-based applications, which exploit the channel sparsity in the delay-Doppler domain, and different complexity-performance tradeoffs have been achieved by modeling the range and range-rate of any path component as either integer or fractional multiples of the range and range-rate resolutions, respectively: [^1] For a given path component, the delay and Doppler parameters are equivalent to its range and range-rate, respectively, with a convenient scaling factor; these terms are often used interchangeably in the literature, and, for the remainder of this study, we will prefer the latter terminology. see [20] and references therein. For example, the work in [8] has derived a widely-used input-output signal model under the assumption of integer range and fractional range-rate for both ideal and rectangular shaping filters and has presented a message passing algorithm for data symbol detection. [^2] Hereafter, integer (fractional) range or range-rate means that the range or range-rate is an integer (fractional) multiple of the respective resolution. Leveraging the model in [8], a pilot-aided channel estimation has been developed in [21]; here, pilot, guard, and data symbols are arranged so as to avoid interference between pilot and data symbols at the receiver, thus facilitating channel estimation and data detection. Another channel estimation method based on a sparse Bayesian learning algorithm has been provided in [22] and, then, extended in [23] to handle both fractional range and range-rate. More recently, a channel estimation method relying on the insertion of random pilots has been presented in [24]. On a different side, leveraging the model in [8], the work in [12] has proposed to exploit existing OTFS communication signals for radar sensing to implement a simplified form of communication-centric ISAC. This idea has been further elaborated in [25], where the Cramér-Rao lower bounds in the estimation of the range and range-rate are derived, in [26], where different pulse shapes are investigated, in [27], where a modified orthogonal matching pursuit algorithm with fractional refinement is proposed, and in [28], where the estimation of the target parameter is formulated as an atomic norm minimization. The 3D-structured sparsity of MIMO-OTFS channel is further explored in [29], where the downlink channel estimation problem is formulated as a sparse signal recovery problem. In addition, OTFS signals have been used for joint localization and communications in [30] and for target tracking in [31].
@@
Previous studies assume that the length (range) of a channel path between the transmitter and receiver remains constant within the OTFS frame, which is typically valid for terrestrial applications. However, this assumption may not hold true for satellite-based space sensing, as the significant relative motion between the transmitter, monitored target, and receiver can cause range migration during the OTFS frame. With a transmitter on a LEO satellite and a radar receiver on an aircraft at a lower altitude, non-cooperative space targets (e.g., space debris, asteroids, spacecrafts, missiles, and unresponsive satellites) may have a bistatic range-rate of several km/s; for a system with $M=512$ subcarriers, $N=128$ symbol intervals, a subcarrier spacing of $\Delta=15$ kHz, a symbol interval of $T=1/\Delta$, and a bistatic range-rate of $\bar{v}=15$ km/s, the bistatic range variation in the OTFS frame is $\bar{v}NT=128$ m and greatly exceeds the bistatic range resolution $c/(M\Delta)\approx 39$ m, where $c$ is the speed of light.
@@
In this study, we investigate the use of OTFS communication signals emitted by LEO satellites for estimating the initial range, range-rate, and amplitude of space targets experiencing range migration in the OTFS frame, henceforth referred to as high-speed targets for brevity. We follow a communication-centric approach where the existing waveform and protocol are not modified to maintain the communication performance unaltered; instead, the focus is on the design of the opportunistic radar receiver and on the analysis of its achievable performance. Sensing high-speed targets is a challenging problem, and previous studies in the radar literature have primarily utilized unmodulated or modulated coherent pulse trains (typically, a train of linear frequency modulated pulses) and developed motion compensation strategies operating in the time or frequency domain to integrate consecutive pulses [32, 33, 34, 35]. However, the impact of range migration on OTFS signals is unexplored, and existing motion compensation methodologies cannot directly be applied in the delay-Doppler domain when off-the-shelf OTFS modulator and demodulator are employed. This study aims to provide initial insights on parameter estimation of high-speed targets using OTFS communication signals, without changing the structure of existing transceivers. The major contributions are summarized below.
@@
  • We derive a novel model for the echo produced by a high-speed target considering both ideal and rectangular shaping filters; in particular, the target range is assumed constant over a block of a few symbols and changes from block-to-block consistently with the range-rate. The proposed model subsumes the one in [8] when the target range migration is negligible and the target range is an integer multiple of the range resolution.
@@
  • We demonstrate that the response of high-speed targets maintains a sparse structure in the delay-Doppler domain, determined entirely by the initial range and range-rate. Additionally, we show that range migration results in a spread of this response across both dimensions.
@@
  • We propose an approximate implementation of the maximum likelihood (ML) estimator for target parameters, which avoids exhaustive two-dimensional search in the delay-Doppler plane. First, a block orthogonal matching pursuit (BOMP) algorithm [36] is used to obtain coarse information on the initial range and range-rate. Next, a refined estimate of the target parameters is derived using a bank of matched filters focused on a narrower delay-Doppler region. The proposed procedure can be combined with the CLEAN algorithm in [37, 38, 39, 40, 16] to handle multiple targets.
  • Numerical simulations are conducted to validate the proposed echo model and assess the performance of the target parameter estimator under both ideal and rectangular shaping filters. Results are also compared against those obtained using the methods in [21] and [29] that assume no range migration occurs within the OTFS frame.
@@
The remainder of the paper is organized as follows. Sec. II presents the system description and the design assumptions. Secs. III and IV provide the proposed signal model for ideal and rectangular shaping filters, respectively. Sec. V illustrates the proposed procedure for target parameter estimation. Sec. VI discusses some numerical examples. Sec. VII contains the conclusions. Finally, some mathematical derivations are deferred to the Appendix.
@@ II System description
Consider a sensing system consisting of a communication transmitter and a radar receiver. The communication transmitter is mounted on a LEO satellite, uses an OTFS modulation, and operates with a carrier frequency $c/\lambda$, where $\lambda$ is the carrier wavelength; the emitted waveform is specified by the underlying communication protocol and cannot be modified to adjust the radar needs. The radar receiver aims to estimate the initial range (delay), range-rate (Doppler shift), and amplitude of high-speed targets illuminated by the communication transmitter; it may be mounted on the same platform hosting the communication transmitter (monostatic configuration) or on a separate device (bistatic configuration), e.g., another LEO satellites, an aircraft, a high altitude platform, or a terrestrial station. The receiver is synchronized with the transmitter and knows the emitted symbols: for a monostatic configuration, this is readily verified [15, 16], whereas, for a bistatic configuration, this requires the presence of some coordination [^3] For example, when the radar receiver is on a different LEO satellite, the coordination may be established via an inter-satellite link. or a separate reference channel able to demodulate the communication signal [41, 42, 43].
For illustration, we consider a bistatic radar configuration, as depicted in Fig. 1. In this setup, the bistatic range of a target is defined as the total distance traveled by the probing signal (first from the communication transmitter to the target, and then from the target to the radar receiver). Correspondingly, the bistatic range-rate refers to the time derivative of the bistatic range, i.e., the rate at which the bistatic range changes due to motion of the transmitter, the target, or the receiver. It is important to note that the following analysis also applies to the monostatic case, where the transmitter and receiver are co-located; in such a configuration, the bistatic range and range-rate simplify to twice the monostatic range and range-rate, respectively. To streamline the exposition, we will omit the adjective "bistatic" from this point forward and use the terms range and range-rate to refer to the bistatic quantities unless otherwise specified. Finally, we begin by assuming the presence of a single target to simplify the initial discussion.
@@ II-A Transmit signal
Consider an OTFS frame with bandwidth $M\Delta$ and duration $NT$, where $M$ and $N$ are the number of subcarriers and symbol intervals, respectively, $\Delta$ is the subcarrier spacing, and $T=1/\Delta$ is the symbol interval, while $1/(M\Delta)$ and $1/(NT)$ are the delay and Doppler resolutions, respectively. Up to a positive scaling factor accounting for the radiated energy, the baseband signal emitted by the communication transmitter is
@@
for $n=0,\ldots,N-1$ and $m=0,\ldots,M-1$. By considering $g_{\rm tx}(t)$ time limited in $[0,T]$, then $x(t)$ has support $[0,NT]$. Also, by assuming the spectrum of $g_{\rm tx}(t)$ contained in $\Delta/2,\Delta/2]$, then the spectrum of $x(t)$ is contained in $\Delta/2,M\Delta-\Delta/2]$. [^4] Since a time limited signal cannot be strictly band limited, some spectral leakage outside $\Delta/2,M\Delta-\Delta/2]$ is present. In practice, the transmit shaping filter may not be exactly zero outside $[0,T]$ to reduce such spectral leakage; in this latter case, $x(t)$ is only approximately contained in $[0,NT]$.
II-B Design assumptions
@@
where $\bar{d}$ and $\bar{v}$ are the initial range and the range-rate, respectively, defined as
@@
Denote by $\bar{r}_{\min}$ and $\bar{r}_{\max}$ the minimum and maximum ranges, respectively, and by $\bar{v}_{\min}$ and $\bar{v}_{\max}$ the minimum and maximum range-rates, respectively. These values are tied to the prior uncertainty on the target location and mobility, as resulting from a previous detection phase (if the radar is operating in search mode) or as predicted by the tracking stage (if the radar is operating in tracking mode), and are assumed known. Also, let $r_{\max}=\bar{r}_{\max}-\bar{r}_{\min}$ and $v_{\max}=\bar{v}_{\max}-\bar{v}_{\min}$ be the length of the inspected range interval and range-rate interval, respectively. Then, we assume
@@
so as to avoid ambiguity in the estimation of the initial range and the range-rate of the target, respectively.
Finally, let $N$ be a positive integer multiple of $B$. Then, we assume that $\bar{r}(t)$ is constant over at least $N/B$ consecutive symbol intervals: we refer to this as the $B$-fold stop-and-go approximation on the range variation over the OTFS frame. Under this design assumption, we can express:
@@
with $T_{B}=NT/B$ and $b=0,\ldots,B-1$. For a fixed system bandwidth, increasing $B$ improves the model’s accuracy, particularly when the target exhibits higher range-rates or the OTFS frame duration is long (see Sec. VI for further discussion). Notably, prior studies assume $B=1$ at the design stage, effectively neglecting any range migration across the entire OTFS frame. In contrast, setting $B=N$ limits this approximation to only a single symbol interval, thereby capturing range migration with a much finer granularity.
@@ II-C Received signal
is the baseband representation of the target echo with respect to the reference frequency $\frac{c+\bar{v}_{\min}}{\lambda}$ and $r(t)=\bar{r}(t)-\bar{r}_{\min}$ is the excess target range (i.e., the target range in excess with respect to $\bar{r}_{\min}$). The support of $y(t)$ is contained in the time interval $0,NT+r_{\max}/c]$, and its spectrum is negligible outside the frequency interval $\Delta/2,M\Delta+v_{\max}/\lambda-\Delta/2]$.
Define the excess initial range $d=\bar{d}-\bar{r}_{\min}\in[0,r_{\max}]$, the excess range-rate $v=\bar{v}-\bar{v}_{\min}\in[0,v_{\max}]$, and $\alpha=\bar{\alpha}\mathrm{e}^{-\mathrm{i}2\pi\frac{d}{\lambda}}$. If the variation of the target range in a symbol interval can be ignored (which is the case under the $B$-fold stop-and-go approximation), then (10) can be expanded as
@@
Notice here that $y_{n}(t)$ is the output of a doubly-selective channel with delay-Doppler spread function $h_{n}(\tau,\nu)$ when the input is $x_{n}(t)$; the spread function depends upon $n$, as the target range can change during the OTFS frame.

### v3

@@ Abstract
This study investigates a communication-centric integrated sensing and communication system that utilizes orthogonal time-frequency space (OTFS) modulated signals emitted by low Earth orbit satellites to estimate the parameters of space targets experiencing range migration, hereinafter referred to as high-speed targets. Leveraging the signal samples produced by off-the-shelf OTFS demodulators, we derive a novel input-output model for the echo generated by a high-speed target when ideal and rectangular shaping filters are employed. Our findings reveal that the target response exhibits a sparse structure in the delay-Doppler domain, whose support is determined by the target initial-range and range-rate. Range migration induces a structured spread of this response, which is explicitly characterized in the paper and differs from that in previous models. We propose an approximate implementation of the maximum likelihood estimator for the target initial-range, range-rate, and amplitude. The estimation process first obtains coarse information on the target response using a block orthogonal matching pursuit algorithm, followed by a refinement step based on a bank of matched filters focused on a smaller initial-range/range-rate region. The proposed single-target procedure is extended to multiple targets via iterative estimation, reconstruction, and cancellation of dominant echoes. Finally, numerical examples are provided to evaluate the estimation performance.
@@ I Introduction
Future wireless networks and Internet-of-Things (IoT) systems will utilize non-terrestrial stations to ensure seamless and ubiquitous connectivity [1]. Among space transceivers, low Earth orbit (LEO) satellites are attracting significant attention due to their low latency, high data rates, and capability to support integrated sensing and communication (ISAC) [2, 3, 4], enabling applications such as environmental monitoring, global navigation, and space exploration.
@@
Orthogonal time frequency space (OTFS) modulation is a promising transmission scheme for future LEO constellations [5]. OTFS-based transceivers exploit channel diversity across both time and frequency, thus providing improved robustness in doubly-dispersive fading channels compared to orthogonal frequency division multiplexing [6, 7]. They are also suitable for radar sensing, as the underlying signal processing resembles that of pulsed-Doppler radars [8]; in particular, the range and range-rate parameters of the detected objects are inferred from the estimated delay-Doppler channel response. [^1] In this work, the range is defined as the propagation distance of the signal path, while the range-rate is defined as its time derivative. Hence, the range and range-rate provide a convenient physical parameterization of the delay and Doppler shifts along the propagation path, respectively. These features enable OTFS-based LEO satellites to support ISAC in various forms [9, 10]. Communication-centric ISAC appears particularly promising in this context, as existing signals can be reused for sensing at low cost without altering the underlying communication protocol [11, 12, 13, 14, 15]; this approach is considered in this study.
@@
A large body of work has addressed signal detection and channel estimation for OTFS systems by exploiting the inherent sparsity of the delay-Doppler representation when range migration over the processed frame can be ignored, an assumption that is usually valid for terrestrial scenarios; see [16] and references therein. For example, [6] derived a widely-used input-output signal model under the assumption of integer range and fractional range-rate for both ideal and rectangular shaping filters, along with a message passing algorithm for data detection. [^2] Integer (fractional) range or range-rate means that the range or range-rate is an integer (fractional) multiple of the respective resolution. Based on this model, pilot-aided channel estimation schemes have been proposed in [17], where guard symbols are introduced to avoid interference between pilot and data symbols at the receiver. Sparse Bayesian learning approaches have been developed in [18] and extended in [19] to handle both fractional range and range-rate. On the sensing side, [11] proposed to exploit OTFS communication signals for radar sensing within a communication-centric ISAC framework. This idea has been further developed in several directions, including the derivation of Cramér–Rao lower bounds for range and range-rate estimation [20], the analysis of pulse shaping effects [21], the design of sparse recovery algorithms with fractional refinement [22], and atomic norm-based formulations for target parameter estimation [23]. The 3D structured sparsity of MIMO-OTFS channels has been investigated in [24], where the downlink channel estimation problem is formulated as a sparse signal recovery task. In addition, OTFS signals have been employed for joint localization and communication in [25] and for target tracking in [26].
@@
Assuming the range of each propagation path to be constant may no longer be valid in satellite-based sensing. For instance, in a scenario with the communication transmitter on a LEO satellite and the radar receiver on an airborne platform, non-cooperative space targets (e.g., debris, asteroids, spacecraft, or missiles) may exhibit range rates of several km/s. When considering a system with $M=512$ subcarriers, $N=128$ symbol intervals, a subcarrier spacing of $\Delta=15$ kHz, a symbol interval of $T=1/\Delta$, and a range rate of $\bar{v}=15$ km/s, the range variation over the OTFS frame is $\bar{v}NT=128$ m, which greatly exceeds the range resolution $c/(M\Delta)\approx 39$ m, where $c$ is the speed of light.
@@
This work investigates the use of OTFS communication signals for estimating the parameters of high-speed targets undergoing range migration within the OTFS frame. We adopt a communication-centric approach in which the waveform is not modified to preserve communication performance, and we focus on characterizing the target echo model, designing the radar receiver, and analyzing estimation performance. Sensing high-speed targets is challenging, and prior radar studies have primarily relied on unmodulated or modulated coherent pulse trains (typically, a train of linear frequency-modulated pulses) and developed motion-compensation strategies in the time or frequency domain to coherently integrate consecutive pulses [27, 28, 29, 30]. However, the impact of range migration on OTFS signals remains unexplored, and existing motion-compensation techniques cannot be directly applied in the delay-Doppler domain under standard modulation and demodulation processing. This study aims to provide initial insights into parameter estimation for high-speed targets using OTFS communication signals, without altering the transceiver structure. The major contributions are summarized below.
@@
  • We derive a model for the echo produced by a high-speed target in the delay-Doppler domain, considering both ideal and rectangular shaping filters; the target range is assumed constant over a block of a few symbols, but changes from block to block consistently with the range-rate. Ideal filters yield a circular convolution model, whereas rectangular filters break this convolution structure because of inter-carrier interference (ICI) and inter-symbol interference (ISI). This model subsumes the one in [6] when the target-range migration is negligible.
@@
  • We demonstrate that the response of high-speed targets maintains a sparse structure in the delay-Doppler domain, determined by its initial-range (i.e., the target range at the beginning of the OTFS frame) and range-rate. Specifically, range migration results in an expanded support of this response not captured by existing models.
@@
  • We propose an approximate implementation of the maximum likelihood (ML) estimator for target parameters, which avoids exhaustive search over the entire parameter space. First, a block orthogonal matching pursuit (BOMP) algorithm [31] is used to obtain coarse information on the initial-range and range-rate. Next, a refined estimate of the target parameters is derived using a bank of matched filters focused on a narrower region. The proposed two-step estimator can be combined with the CLEAN algorithm in [32, 33, 34, 35, 12] to handle multiple targets.
  • Numerical simulations are conducted to validate the proposed echo model and assess the performance of the proposed two-step estimator under different system configurations and numbers of targets, also in comparison with the estimators in [17] and [24].
@@
The remainder of the paper is organized as follows. Sec. II presents the system description. Sec. III presents the delay–Doppler input–output model under ideal and rectangular shaping filters. Sec. IV contains the proposed procedure for target parameter estimation. Sec. V discusses some numerical examples. Sec. VI contains the conclusions. Finally, the appendices contain some mathematical derivations.
@@ II System description
Consider a system consisting of a communication transmitter and a radar receiver. The communication transmitter is mounted on a LEO satellite, uses an OTFS modulation, and operates at a carrier frequency $c/\lambda$, where $\lambda$ is the carrier wavelength; the emitted waveform is specified by the underlying communication protocol and is not modified here. The radar receiver aims to estimate the initial-range, range-rate, and amplitude of high-speed targets illuminated by the communication transmitter; it may be mounted on the same platform hosting the communication transmitter (monostatic configuration) or on a separate device (bistatic configuration), e.g., another LEO satellite, an aircraft, a high altitude platform, or a terrestrial station. The receiver is synchronized with the transmitter and knows the emitted symbols: for a monostatic configuration, this is readily verified [12], whereas, for a bistatic configuration, this requires some coordination [^3] When the radar receiver is on a different LEO satellite, the coordination may be established via an inter-satellite link. or a separate reference channel [36, 37, 38]. For illustration, in this study we consider the bistatic configuration depicted in Fig. 1.
We assume point-like targets, each generating a single line-of-sight propagation path. The point-like model is appropriate when the target extent is small relative to the range resolution and allows us to isolate and analyze the effects of range migration. The single-path model is consistent with the considered space-sensing scenario, where secondary scattering components are typically negligible or absent. Extensions to distributed targets or to more general propagation conditions involving multiple echoes from the same target are left for future work. For clarity of exposition, we initially focus on a single target, while multiple targets are addressed in Sec. IV-B.
@@ II-A Transmit signal
Consider an OTFS frame with bandwidth $M\Delta$ and duration $NT$, where $M$ and $N$ are the number of subcarriers and symbol intervals, respectively, $\Delta$ is the subcarrier spacing, and $T=1/\Delta$ is the symbol interval. Accordingly, the corresponding range and range-rate resolutions are [36]
$R_{\rm r}$ $=c/(M\Delta),$ (1a)
$R_{\rm rr}$ $=\lambda/(NT),$ (1b)
[…]
@@
for $n=0,\ldots,N-1$ and $m=0,\ldots,M-1$. By considering $g_{\rm tx}(t)$ time limited in $[0,T]$, then $x(t)$ has support $[0,NT]$. Also, by assuming the spectrum of $g_{\rm tx}(t)$ is contained in $0,\Delta]$, then the spectrum of $x(t)$ is contained in [^5] Since a time-limited signal cannot be strictly band-limited, some spectral leakage outside $0,M\Delta]$ is present. In practice, the transmit shaping filter may not be exactly zero outside $[0,T]$ to reduce such spectral leakage; in this latter case, $x(t)$ is only approximately contained in $[0,NT]$. $0,M\Delta]$.
II-B Target kinematics and design assumptions
@@
where $\bar{d}$ and $\bar{v}$ are the initial-range and the range-rate, respectively, defined as
@@
Hereafter, we make the following design assumptions.
II-B1 Unambiguous initial-range and range-rate intervals
Denote by $\bar{r}_{\min}$ and $\bar{r}_{\max}$ the minimum and maximum initial-ranges, respectively, and by $\bar{v}_{\min}$ and $\bar{v}_{\max}$ the minimum and maximum range-rates, respectively. These values are tied to the prior uncertainty on the target location and mobility, as resulting from a previous detection phase (if the radar is operating in search mode) or as predicted by the tracking stage (if the radar is operating in tracking mode), and are assumed known. Also, let $r_{\max}=\bar{r}_{\max}-\bar{r}_{\min}$ and $v_{\max}=\bar{v}_{\max}-\bar{v}_{\min}$ be the length of the inspected initial-range and range-rate intervals, respectively. Then, we assume
@@ II-B1 Unambiguous initial-range and range-rate intervals
so as to avoid ambiguity in the estimation of the initial-range and the range-rate of the target, respectively. [^6] For the parameters adopted in Sec. V ($\Delta=15$ kHz, $M=512$, $N=128$, carrier frequency $4$ GHz), we have $c/\Delta\approx 19.9$ km and $\lambda\Delta\approx 1.12$ km/s, respectively. By varying $\bar{r}_{\min}$ and $\bar{v}_{\min}$, any initial-range interval of length $r_{\max}<19.9$ km and any range-rate interval of length $v_{\max}<1.12$ km/s can be jointly inspected without ambiguity.
II-B2 Blockwise constant-range approximation
Let $B$ be a divisor of $N$. Then, the OTFS frame can be partitioned into $B$ equal blocks of $N/B$ symbol intervals. For receiver design, the target range is assumed constant within each block. Accordingly, we write
@@ II-B2 Blockwise constant-range approximation
where $T_{B}=NT/B$ and $b=0,\ldots,B-1$. Increasing $B$ improves the model accuracy, particularly when the target exhibits higher range-rates or when the OTFS frame duration is long (see Sec. V for further discussion). A practical design guideline is to select $B$ such that the range variation over each block, i.e., $\bar{v}_{\max}NT/B$, is significantly smaller than the range resolution $R_{\rm r}$ (see also Sec. V). Prior studies assume $B=1$, which is appropriate only for applications in which target-range migration within the OTFS frame can be neglected.
@@ II-C Received signal
is the baseband representation of the target echo with respect to the reference frequency $\frac{c+\bar{v}_{\min}}{\lambda}$ and $r(t)=\bar{r}(t)-\bar{r}_{\min}$ is the excess target range (i.e., the target range in excess with respect to $\bar{r}_{\min}$). Note that the support of $y(t)$ is contained in the time interval $0,NT+r_{\max}/c]$, and its spectrum is negligible outside the frequency interval $0,M\Delta+v_{\max}/\lambda]$.
Define the excess initial-range $d=\bar{d}-\bar{r}_{\min}\in[0,r_{\max}]$, the excess range-rate $v=\bar{v}-\bar{v}_{\min}\in[0,v_{\max}]$, and $\alpha=\bar{\alpha}\mathrm{e}^{-\mathrm{i}2\pi\frac{d}{\lambda}}$. Ignoring the variation of the target range in a symbol interval, then (11) can be expanded as
@@
Notice that $y_{n}(t)$ is the output of a doubly-selective channel with delay-Doppler spread function $h_{n}(\tau,\nu)$ when the input is $x_{n}(t)$; the spread function depends upon $n$, as the target range can change during the OTFS frame.